# Continuous Temporal Domain Generalization

### v1

@@ 5 Experiment
In this section, we present the performance of the Koodos in comparison to other approaches through both quantitative and qualitative analyses. Our evaluation aims at 1) assessing the generalizability of Koodos in continuous temporal domains; 2) assessing whether Koodos captures the correct underlying data dynamics; and 3) assessing whether Koodos can use inductive bias to guide the behavior of predictive models. Detailed experiment settings and results (e.g., hyperparameter setting, sensitivity analysis, ablation study, and scalability analysis) are demonstrated in the Appendix A.1. Besides, since the TDG is a special case of the CTDG, we also conducted experiments on the traditional discrete temporal domain data. Results are shown in Appendix B.
Datasets. We compare with classification datasets: Rotated Moons (2-Moons), Rotated MNIST (Rot-MNIST), Twitter Influenza Risk (Influenza), and Yearbook; and the regression datasets: Tropical Cyclone Intensity (Cyclone), House Prices (House). More details can be found in Appendix A.1.1.
Baselines. We employ three categories of baselines: Practical baselines, including 1) Offline; 2) LastDomain; 3) IncFinetune. Discrete temporal domain generalization methods, including 1) DRAIN [1]; 2) DRAIN-$\Delta t$. Continuous temporal domain generalization methods, including 1) DeepODE [7]. Comparison method details can be found in Appendix A.1.2.
Metrics. Error rate (%) is used for classification tasks. As the Influenza dataset has imbalanced labels, the Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) curve is used. Mean Absolute Error (MAE) is used for regression tasks. All models were trained on training domains and then deployed on all unseen test domains. Each method’s experiments were repeated five times, with mean results and standard deviations reported. Detailed parameter settings for each dataset are provided in Appendix A.1.3.
Table 1: Performance comparison on continuous temporal domain datasets. The classification tasks report Error rates (%) except for the AUC for the Influenza. The regression tasks report MAE.
@@
2-Moons Rot-MNIST Influenza Yearbook Cyclone House
@@ 5.1 Quantitative Analysis: Generalization Across Continuous Temporal Domains
We first illustrate the performance of our proposed method against baseline methods, highlighting results from Table 1. Koodos demonstrates exceptional generalizability across continuous temporal domains, significantly outperforming the second-best methods on datasets such as 2-Moons, Influenza, and Cyclone. We observe that all baselines struggle with synthetic datasets, particularly challenged by the continuous, significant concept shifts (i.e., 18 degrees of rotation per second). In real-world datasets, while DRAIN shows some effectiveness on the House dataset, its performance across other datasets underscores the limitations of using probabilistic discrete methods (i.e., LSTM) for continuous domain modeling. Despite attempts to adjust for temporal irregularities, neither DRAIN-$\Delta t$ nor DeepODE achieves satisfactory results, as they never attempt to synchronize the data and model dynamics. In contrast, Koodos establishes a promising approach and a benchmark in CTDG tasks, with quantitative analysis firmly confirming our approach’s superiority.
@@
Figure 3: Visualization of Decision Boundary of the 2-Moons Dataset (purple and yellow colors indicate the different data class area and the red line identifies the decision boundary) From top to bottom: demonstrate two baseline methods compared with ours; From left to right: demonstrate selected of the test domains (red docs); All models are learned using data before the last train domain.
Figure 4: Interpolated and Extrapolated Predictive Model Trajectories
Figure 5: Eigenvalue Distribution of the Koopman Operator
5.2 Qualitative Analysis: Model Dynamic with Data Dynamics
@@
We conducted a qualitative comparison of different models by visualizing their decision boundaries on the 2-Moons dataset, as depicted in Fig. 3. Each row represents a different method: DRAIN-$\Delta t$, DeepODE, and Koodos, with the timeline at the bottom tracking progress through the test domains. DRAIN-$\Delta t$ displays the highest error rate, showing substantial deviation from the anticipated trajectories, especially after the third domain. We also observe that DRAIN-$\Delta t$ seems to freeze when predicting multiple steps, likely due to its underlying model, DRAIN, uses a recursive training strategy within a single domain and is explicitly designed for one-step prediction. DeepODE shows a relatively better performance. It benefits from leveraging ODEs to maintain continuity in the model dynamics. However, the nonlinear variation of the predictive model parameters complicates its ability to abstract and simplify the real dynamics. Its predictions start close to the desired paths but diverge over time. Finally, Koodos exhibits the highest performance with clear and concise boundaries, consistently aligning with the actual dynamics and maintaining high fidelity across all tested domains, showcasing its robustness in continuous temporal domain modeling and generalization.
@@
Fig. 5 demonstrates the space-time evolution of the generalized predictive model from Koodos. By applying t-SNE, the predictive model parameters are reduced to a 2-dimensional representation space, plotted against the time on the Z-axis. We interpolated the model states (35 seconds displayed in blue-yellow line) among the training domain states (marked by blue docs) in steps of 0.2 seconds, and similarly extrapolated 75 steps (15 seconds displayed in red line). This visualization clearly shows that the model dynamics are synchronized with the data dynamics: the interpolation creates a cohesive, upward-spiraling trajectory transitioning from the first to the last training domain, while the extrapolation correctly extends this trajectory outward into a new area, demonstrating the effective of Koodos from another intuitive way. We show the space-time evolution of baseline models in Fig. 8.
5.3 Analysis and Control of Predictive Models
@@
The learned Koopman operator provides valuable insights into the dynamics of generalized predictive models. We analyze the behavior of the learned Koodos model on the 2-Moons dataset, focusing on the eigenvalues of the Koopman operator. As illustrated in Fig. 5(a), the eigenvalues are distributed across both stable (Re<0) and unstable (Re>0) regions on the complex plane. The spectral distribution suggests that while Koodos performs effectively across all tested domains, it will eventually diverge towards instability and finally collapse. To validate, we extended the extrapolation of the Koodos model to an extremely long-term (i.e., 35 seconds future). Results depicted in Fig. 9(a) demonstrate that the predictive model’s trajectory significantly deviates from the anticipated spiral path, suggesting that the prediction of Koodos over such long-term will end in error accumulation.
@@
Koodos’s innovative design allows it to incorporate knowledge that extends beyond the observable data. By configuring the Koopman operator as $\mathcal{K}=B-B^{T}$, we ensure that all eigenvalues of the final learned $\mathcal{K}$ are purely imaginary (termed Koodos+), promoting stable and periodic behavior. This adjustment is reflected in Fig. 5(b), where the eigenvalues are tightly clustered around the imaginary axis. As shown in Fig. 9(b), the embeddings and trajectories of Koodos+ are cohesive and maintain stability over extended periods. Remarkably, even with 85 seconds of extrapolation, Koodos+ shows no signs of performance degradation, highlighting the significance of human inductive bias in improving the robustness and reliability of the generalized model.
@@ A.1.1 Dataset Details
    Influenza The Influenza dataset [51] utilizes tweets data to predict flu intensity. We randomly collected streaming tweets starting at 50 arbitrary timestamps lasting 7 days during the 2010-2014 flu seasons and then examined the volume of disease-related terms to predict current flu trends. The result is validated against the Influenza-Like Illness (ILI) activity levels reported by the Centers for Disease Control and Prevention (CDC). The continuous concept drift in this dataset is characterized by the fluctuations in tweet volumes over the years and the variation in flu activity throughout the season.
@@
    Yearbook The Yearbook dataset [48] consists of frontal-facing yearbook portraits collected between 1930 and 2013 from 128 American high schools. The task is to classify gender from face images. We randomly selected 40 different years of data from the 84-year dataset to simulate missing domain data collection in a realistic situation. This means that the time difference between two consecutive domains is not necessarily 1 year. The Yearbook dataset provides a visual record of evolving fashion styles and social norms across the decades.
@@
    Cyclone The Cyclone dataset [6] is collected through satellite remote sensing and is designed for the task of tropical cyclone image to wind intensity estimation. This dataset tracks each cyclone throughout its entire lifecycle by capturing satellite photographs every three hours. We focused on cyclone data from the West Pacific region covering the years 2014 to 2016 and formed 72 continuous domains. The datasets are event-triggered, and each domain is indexed by the date the cyclone occurred. The variation in wind strength associated with the seasonal dates results in a continuous concept drift.
@@
To clearly characterize the temporal arbitrariness of the continuous temporal domains, we plot the time distribution of each dataset in Fig. 6, where the blue line represents the timestamp of one specific domain. We use the last 30% domain of each dataset as test domains, which are shaded in gray.
Figure 6: Distribution of Domains in Time for Continuous Temporal Domain Datasets
@@ A.1.2 Comparison Methods
    Practical Baseline (1) Offline Model: This model operates without temporal variations and is trained using Empirical Risk Minimization (ERM) across all source domains. (2) LastDomain: This model is also trained without temporal variations but focuses solely on the last source domain using ERM. (3) IncFinetune: This approach begins by applying the baseline method at the initial time point and subsequently fine-tunes the model at each following time point using a lower learning rate. This incremental finetuning strategy biases the training towards more recent data.
@@
    Discrete Temporal Domain Generalization Baseline Continuous temporal domain tasks require consideration of domains for the long future rather than a limited number of steps. Existing methods are mainly unable to deal with this problem. Among them, although DRAIN [1] is designed only for predicting one-step future prediction, its internal LSTM structure has the potential of long-term future prediction, so we implemented DRAIN as a discrete baseline. We implemented two versions of it: (1) DRAIN: This approach tackles the temporal domain generalization problem by proposing a dynamic neural network, which uses an LSTM to capture the evolving pattern of a neural network, during the test, DRAIN will consider all future test domains as one domain, and utilizes the LSTM to predict the future parameters of this target domain. (2) DRAIN-$\Delta t$: A variant of DRAIN, in which the internal LSTM is changed to a continues recurrent units [38] to model irregular distributed domains. Under this setting, DRAIN-$\Delta t$ treat the test domians as a irregular time series and predict the model state at that given moment.
@@ A.1.3 Model Configuration
We use autoencoder to achieve the $\varphi$ and $\varphi^{-1}$, which we called Encoder and Decoder in the following. A one-layer linear transformation layer (named Dynamic Model) without bias is served as the Koopman operator. In some complex data dynamics, the Koopman operator requires increasing the operator dimension (up to infinite dimensions) to obtain a better approximation of the dynamics. Since we cannot grow the dimension indefinitely, alternatively, we use two linear layers with one ReLU to approximate the Koopman operator with infinite dimensional width. In practice we found that two linear layers work well enough. Additionally, a Predictive Model for each domain is trained to provide initial parameter values for the Dynamic Model, and to facilitate the computation of loss as detailed in Section 4.3. We use NeuralODEs [7] as our ODEs solvers. For large models with massive parameters, we treat the head network as a feature extractor shared by all domains, inferring only the state of the remaining layers.
We use Adam Optimizer for all our experiments. All experiments are conducted on a 64-bit machine with two 20-core Intel Xeon Silver 4210R CPU @ 2.40GHz, 378GB memory, and four NVIDIA GeForce RTX 3090. We set hyperparameters for each comparison method with respect to the recommendation in their original paper, and we specify the architecture as well as other details for each dataset’s experiments as follows.
@@
    Rotated 2-Moons The Predictive Model consists of 2 hidden layers, with a dimension of 50 each. We use ReLU layer after each layer and a Sigmoid layer after the output layer. The Encoder and Decoder are both a 4-layer Multi-Layer Perceptron (MLP), with dimensions of $[1024,512,128,32]$ for each layer. A 32-dimensional linear layer Dynamic Model serves as the Koopman operator. The learning rate for the Predictive Model is set at $1\times 10^{-2}$, while for the others is set at $1\times 10^{-3}$.
@@
    Rotated MNIST The Prediction Model features a convolutional neural network architecture comprising three convolutional layers with channel $[32,32,64]$ and a kernel size of 3. Each convolutional layer is followed by a ReLU activation layer and a max pooling layer with a kernel size of 2. Following the flattening, the network includes two hidden linear layers with dimensions $[128,10]$. We treat the convolutional layer as the feature extractor. The Encoder and Decoder are both a 4-layer MLP, with dimensions of $[1024,512,128,32]$ for each layer. A 32-dimensional 2 linear layers Dynamic Model serves as the Koopman operator. The learning rate for all parts is set at $1\times 10^{-3}$.
@@
    Influenza The Prediction Model consists of 3 hidden layers, with a dimension of $[128,32,1]$. We use ReLU layer after each layer and a Sigmoid layer after the output layer. The Encoder and Decoder are both a 4-layer MLP, with dimensions of $[1024,512,128,32]$ for each layer. A 32-dimensional 2 linear layers Dynamic Model serves as the Koopman operator. The learning rate for all parts is set at $1\times 10^{-3}$.
@@
    Yearbook The Prediction Model features a convolutional neural network architecture comprising three convolutional layers with channel $[32,32,64]$ and a kernel size of 3. Each convolutional layer is followed by a ReLU activation layer and a max pooling layer with a kernel size of 2. Following the flattening, the network includes three hidden linear layers with dimensions $[128,32,1]$. We treat the convolutional layer as the feature extractor. The Encoder and Decoder are both a 4-layer MLP, with dimensions of $[1024,512,128,32]$ for each layer. A 32-dimensional 2 linear layers Dynamic Model serves as the Koopman operator. The learning rate for all parts is set at $1\times 10^{-3}$.
@@
    Cyclone The Prediction Model features a convolutional neural network architecture comprising four convolutional layers with channel $[32,32,64,64]$ and a kernel size of 3. Each convolutional layer is followed by a ReLU activation layer and a max pooling layer with a kernel size of 2. Following the flattening, the network includes three hidden linear layers with dimensions $[128,32,1]$. We treat the convolutional layer as the feature extractor. The Encoder and Decoder are both a 4-layer MLP, with dimensions of $[1024,512,128,32]$ for each layer. A 32-dimensional 2 linear layers Dynamic Model serves as the Koopman operator. The learning rate for all parts is set at $1\times 10^{-3}$.
@@
    House The Prediction Model consists of 2 hidden layers with a dimension of 400. We use the ReLU layer after each layer. The Encoder and Decoder are both a 4-layer MLP, with dimensions of $[1024,512,128,32]$ for each layer. A 32-dimensional 2 linear layers Dynamic Model serves as the Koopman operator. The learning rate for all parts is set at $1\times 10^{-3}$.

### v2

@@ 5 Experiment
In this section, we present the performance of the Koodos in comparison to other approaches through both quantitative and qualitative analyses. Our evaluation aims at 1) assessing the generalizability of Koodos in continuous temporal domains; 2) assessing whether Koodos captures the correct underlying data dynamics; and 3) assessing whether Koodos can use inductive bias to guide the behavior of the generalization. Detailed experiment settings (i.e., dataset details, baseline details, hyperparameter settings, ablation study, scalability analysis, and sensitivity analysis) are demonstrated in Appendix A.1. Besides, since the TDG is a special case of the CTDG, we also conducted experiments on the traditional discrete temporal domain generalization task. Results are shown in Appendix B.
Datasets. We compare with classification datasets: Rotated Moons (2-Moons), Rotated MNIST (Rot-MNIST), Twitter Influenza Risk (Twitter), and Yearbook; and the regression datasets: Tropical Cyclone Intensity (Cyclone), House Prices (House). More details can be found in Appendix A.1.1.
Baselines. We employ three categories of baselines: Practical baselines, including 1) Offline; 2) LastDomain; 3) IncFinetune; 4) IRM [1]; 5) V-REx [25]. Discrete temporal domain generalization methods, including 1) CIDA [48]; 2) TKNets [57]; 3) DRAIN [2]; 4) DRAIN-$\Delta t$. Continuous temporal domain generalization methods, including 1) DeepODE [11]. Comparison method details can be found in Appendix A.1.2.
Metrics. Error rate (%) is used for classification tasks. As the Twitter dataset has imbalanced labels, the Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) curve is used. Mean Absolute Error (MAE) is used for regression tasks. All models were trained on training domains and then deployed on all unseen test domains. Each method’s experiments were repeated five times, with mean results and standard deviations reported. Detailed parameter settings for each dataset are provided in Appendix A.1.3.
Table 1: Performance comparison on continuous temporal domain datasets. The classification tasks report Error rates (%) except for the AUC for the Twitter dataset. The regression tasks report MAE. ’N/A’ implies that the method does not support the task.
@@
2-Moons Rot-MNIST Twitter Yearbook Cyclone House
@@ 5.1 Quantitative analysis: generalization across continuous temporal domains
We first present the performance of our proposed method against baseline methods, highlighting results from Table 1. Koodos exhibits outstanding generalizability across continuous temporal domains. A key observation is that all baseline models struggle to handle synthetic datasets, particularly challenged by the continuous and substantial concept drift (i.e., 18 degrees of rotation per second). In real-world datasets, methods like CIDA, DRAIN, and DeepODE demonstrate effectiveness in certain cases. However, the performance gap between them and Koodos highlights the importance of explicitly considering continuous temporal modeling. For instance, while DRAIN attempts to address domain dynamics through a probabilistic sequential approach via LSTM, this introduces considerable errors due to the inherent discrete temporal. Moreover, while DRAIN-$\Delta t$ and DeepODE adjust for temporal irregularities accordingly, they fail to adequately synchronize the data and model dynamics, leading to unsatisfactory results. In contrast, Koodos establishes a promising approach and a benchmark in CTDG tasks, with quantitative analysis firmly confirming the approach’s superiority.
@@
Figure 3: Visualization of decision boundary of the 2-Moons dataset (purple and yellow show data regions, red line shows the decision boundary). Top to bottom compares two baseline methods with ours; left to right shows partial test domains (all test domains are marked with red points on the timeline). All models are learned using data before the last train domain.
Figure 4: Interpolated and extrapolated predictive model trajectories. Left: Koodos captures the essence of generalization through the harmonious synchronization of model and data dynamics; Middle: DRAIN, as a probabilistic model, fails to capture continuous dynamics, which is presented as jumps from one random state to another. Right: DeepODE demonstrates a certain degree of continuity, but the dynamics are incorrect.
5.2 Qualitative analysis: data dynamics and the learned model dynamics
@@
We conducted a qualitative comparison of different models by visualizing their decision boundaries on the 2-Moons dataset, as depicted in Fig. 3. Each row represents a different method: DRAIN-$\Delta t$, DeepODE, and Koodos, with the timeline at the bottom tracking progress through test domains. DRAIN-$\Delta t$ displays the highest error rate, showing substantial deviation from the anticipated trajectories, especially after the third domain. We also observe that DRAIN-$\Delta t$ seems to freeze when predicting multiple steps, likely due to its underlying model, DRAIN, uses a recursive training strategy within a single domain and is explicitly designed for one-step prediction. DeepODE shows a relatively better performance. It benefits from leveraging ODEs to maintain continuity in the model dynamics. However, the nonlinear variation of the predictive model parameters complicates its ability to abstract and simplify the real dynamics. Its predictions start close to the desired paths but diverge over time. Finally, Koodos exhibits the highest performance with clear and concise boundaries, consistently aligning with the actual dynamics and maintaining high fidelity across all tested domains, showcasing its robustness in continuous temporal domain modeling and generalization.
@@
Fig. 4(a) demonstrates the space-time evolution of the generalization process of Koodos. By applying t-SNE, the predictive model parameters are reduced to a 2-dimensional representation space, plotted against the time on the Z-axis. We used Koodos to interpolate the model states (35 seconds displayed in blue-yellow line) among the training domain states (marked by blue docs) in steps of 0.2 seconds, and similarly extrapolated 75 steps (15 seconds displayed in red line). The visualization clearly shows that Koodos synchronizes the model dynamics with the data dynamics: the interpolation creates a cohesive, upward-spiraling trajectory transitioning from the first to the last training domain, while the extrapolation correctly extends this trajectory outward into a new area, demonstrating the effective of Koodos from another intuitive way. We also show the space-time evolution of baseline models in Fig. 4(b,c), in which we do not find meaningful patterns of the generalization process.
Figure 5: Eigenvalue distribution of the Koopman operator. Left: $\mathcal{K}$ as learnable; Right: $\mathcal{K}=B-B^{T}$ with $B$ as learnable.
Figure 6: Extremely long-term extrapolated predictive model trajectories in uncontrolled and controlled settings. Left: $\mathcal{K}$ as learnable; Right: $\mathcal{K}=B-B^{T}$ with $B$ as learnable.
5.3 Analysis and control of the generalization process
@@
The learned Koopman operator provides valuable insights into the dynamics of generalized predictive models. We analyze the behavior of the learned Koodos model on the 2-Moons dataset, focusing on the eigenvalues of the Koopman operator. As illustrated in Fig. 5(a), the eigenvalues are distributed across both stable (Re<0) and unstable (Re>0) regions on the complex plane. The spectral distribution suggests that while Koodos performs effectively across all tested domains, it will eventually diverge towards instability and finally collapse. To validate, we extended the extrapolation of the Koodos to an extremely long-term (i.e., 35 seconds future). Results depicted in Fig. 6(a) demonstrate that the generalized model’s trajectory significantly deviates from the anticipated spiral path, suggesting that extremely long-term generalization will end up with the accumulation of errors.
@@
Fortunately, Koodos’s innovative design allows it to incorporate knowledge that extends beyond the observable data. By configuring the Koopman operator as $\mathcal{K}=B-B^{T}$, we ensure that all eigenvalues of the final learned $\mathcal{K}$ are purely imaginary (termed Koodos+), promoting stable and periodic behavior. This adjustment is reflected in Fig. 5(b), where the eigenvalues are tightly clustered around the imaginary axis. As shown in Fig. 6(b), the embeddings and trajectories of Koodos+ are cohesive and maintain stability over extended periods. Remarkably, even with 85 seconds of extrapolation, Koodos+ shows no signs of performance degradation, highlighting the significance of human inductive bias in improving the robustness and reliability of the generalization process.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported by Japan Science and Technology Agency (JST) SPRING, Grant Number JPMJSP2108, the SPRING GX Overseas Dispatch Program, and the Shibasaki Scholarship Foundation.
@@ A.1.1 Dataset details
    Twitter The Twitter dataset [58, 3] utilizes tweet data to predict flu intensity. We randomly collected streaming tweets starting at 50 arbitrary timestamps lasting 7 days during the 2010-2014 flu seasons and then examined the volume of disease-related terms to predict current flu trends. The result is validated against the Influenza-Like Illness (ILI) activity levels reported by the Centers for Disease Control and Prevention (CDC). The continuous concept drift in this dataset is characterized by the fluctuations in tweet volumes over the years and the variation in flu activity throughout the season.
@@
    Yearbook The Yearbook dataset [54] consists of frontal-facing yearbook portraits collected between 1930 and 2013 from 128 high schools. The task is to classify gender from images. We randomly sampled 40 years of data from the 84-year dataset, with each year representing a domain to represent the incomplete temporal domain collection process, which resulted in variable time intervals between consecutive domains. The Yearbook dataset provides a visual record of evolving fashion styles and social norms across the decades.
@@
    Cyclone The Cyclone dataset [10] is collected by satellite remote sensing and is dedicated to the task of tropical cyclone imagery to wind intensity. When each cyclone occurs, the satellite collects a series of images for its entire life cycle as a domain, with the date of its occurrence representing a temporal domain time. We focused on cyclone data from the West Pacific region covering 2014 to 2016 and formed 72 continuous domains. The dataset is event-triggered, and the variation in wind strength associated with the seasonal dates results in a continuous concept drift.
@@
To more clearly illustrate the temporal arbitrariness of the continuous temporal domain, we plot the time distribution of all domains for each dataset in Fig. 7, where the blue line represents the timestamp of one specific domain. We use the last 30% domain of each dataset as test domains, which are marked with gray shading.
Figure 7: The distribution of continuous temporal domain datasets along the time dimension. The test domains are marked with gray shading.
@@ A.1.2 Comparison methods
    Practical Baseline (1) Offline Model: This model operates without temporal variations and is trained using Empirical Risk Minimization (ERM) across all source domains. (2) LastDomain: This model is also trained without temporal variations but focuses solely on the last source domain using ERM. (3) IncFinetune: This approach begins by applying the baseline method at the initial time point and subsequently fine-tunes the model at each following time point using a lower learning rate. (4) IRM: IRM is to train a predictive model over all distributions that do not rely on spurious correlations as much as possible. (5) V-REx: Like IRM, V-REx performs causal identification, while also providing some robustness to changes in the covariate shift by risk extrapolation.
@@
    Discrete Temporal Domain Generalization Baseline Continuous temporal domain tasks require consideration of domains for the long future rather than a limited number of steps. Existing methods are mainly unable to deal with this problem. Among them, although DRAIN [2] is designed only for predicting one-step future prediction, its internal LSTM structure has the potential of long-term future prediction, so we implemented DRAIN as a discrete baseline. We implemented two versions of it: (1) DRAIN: This approach tackles the temporal domain generalization problem by proposing a dynamic neural network, which uses an LSTM to capture the evolving pattern of a neural network, during the test, DRAIN will consider all future test domains as one domain, and utilizes the LSTM to predict the future parameters of this target domain. (2) DRAIN-$\Delta t$: A variant of DRAIN, in which the internal LSTM is changed to a continues recurrent units [44] to model irregular distributed domains. Under this setting, DRAIN-$\Delta t$ treats the test domains as an irregular time series and predicts the model state at that given moment. We also added CIDA [48] and TKNets [57] as comparisons. (3) CIDA treats the domain index as a continuous variable to help the discriminator using a distance-based loss. (4) TKNets trains a prediction model for the target domain by leveraging the evolving pattern among domains.
@@ A.1.3 Model configuration
We use autoencoder to achieve the $\varphi$ and $\varphi^{-1}$, which we call Encoder and Decoder in the following. A linear transformation layer without bias (named Dynamic Model) is served as the Koopman operator. In certain complex data dynamics, the Koopman operator needs to expand its dimensions (potentially up to infinity) for a better approximation of the dynamics. Since it’s not feasible to increase the dimensions endlessly, we approximate the infinitely wide Koopman operator using multiple linear layers with ReLU. In practice, we found that two linear layers work well enough. Additionally, Predictive Models for each domain are trained to provide initial parameter values for Generalized Models. We use NeuralODEs [11] as the ODEs solvers. For large models with massive parameters, we treat the head network as a feature extractor shared by all domains, inferring only the state of the remaining layers. All experiments are conducted on a 64-bit machine with two 20-core Intel Xeon Silver 4210R CPU @ 2.40GHz, 378GB memory, and four NVIDIA GeForce RTX 3090. We use Adam Optimizer for all experiments, and we specify the architecture as well as other details for each dataset as follows:
@@
    Rotated 2-Moons The Predictive Model consists of 3 hidden layers, with a dimension of 50 each. We use a ReLU layer after each layer and a Sigmoid layer after the output layer. The Encoder and Decoder are both a 4-layer Multi-Layer Perceptron (MLP), with dimensions of $[1024,512,128,32]$ for each layer. A 32-dimensional linear layer Dynamic Model serves as the Koopman operator. The learning rate for the Predictive Model is set at $1\times 10^{-2}$, while for the others is set at $1\times 10^{-3}$.
@@
    Rotated MNIST The Prediction Model features a convolutional neural network architecture comprising three convolutional layers with channel $[32,32,64]$ and a kernel size of 3. Each convolutional layer is followed by a ReLU layer and a max pooling layer with a kernel size of 2. Following the flattening, the network includes two linear layers with dimensions $[128,10]$. A dropout layer is added between the linear layers to prevent overfitting. We treat the convolutional layer as the feature extractor. The Encoder and Decoder are both a 4-layer MLP with dimensions of $[1024,512,128,32]$. A 32-dimensional two linear layers Dynamic Model serves as the Koopman operator. The learning rate for all parts is set at $1\times 10^{-3}$.
@@
    Twitter The Prediction Model consists of 3 hidden layers, with a dimension of $[128,32,1]$. We use the ReLU layer after each layer and a Sigmoid layer after the output layer. The Encoder and Decoder are both a 4-layer MLP, with dimensions of $[1024,512,128,32]$ for each layer. A 32-dimensional two linear layers Dynamic Model serves as the Koopman operator. The learning rate for all parts is set at $1\times 10^{-3}$.
@@
    Yearbook The Prediction Model features a convolutional neural network architecture comprising three convolutional layers with channel $[32,32,64]$ and a kernel size of 3. Each convolutional layer is followed by a ReLU layer and a max pooling layer with a kernel size of 2. Following the flattening, the network includes three hidden linear layers with dimensions $[128,32,1]$. A dropout layer is added between the linear layers to prevent overfitting. The convolutional layer is used as the feature extractor. The Encoder and Decoder are both a 4-layer MLP with dimensions of $[1024,512,128,32]$. A 32-dimensional two linear layers Dynamic Model serves as the Koopman operator. The learning rate is set at $1\times 10^{-3}$.
@@
    Cyclone The Prediction Model features a convolutional neural network architecture comprising four convolutional layers with channel $[32,32,64,64]$ and a kernel size of 3. Each convolutional layer is followed by a ReLU layer and a max pooling layer with a kernel size of 2. Following the flattening, the network includes three hidden linear layers with dimensions $[128,32,1]$. We treat the convolutional layer as the feature extractor. The Encoder and Decoder are both a 4-layer MLP, with dimensions of $[1024,512,128,32]$ for each layer. A 32-dimensional 2 linear layers Dynamic Model serves as the Koopman operator. The learning rate for all parts is set at $1\times 10^{-3}$.
@@
    House The Prediction Model consists of 3 hidden layers with a dimension of 400. We use the ReLU layer after each layer. The Encoder and Decoder are both a 4-layer MLP with dimensions of $[1024,512,128,32]$. A 32-dimensional two linear layers Dynamic Model serves as the Koopman operator. The learning rate for all parts is set at $1\times 10^{-3}$.